OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that contains the following points? (–2, 18), (0, 2), (4, 42) y = –2x^2 – 2x – 3 y = –3x^2 + 2x – 2 y = 3x^2 – 2x + 2 y = –2x^2 + 3x + 2

OpenStudy (amistre64):

when x=0, we can narrow down the options to 2 of them

OpenStudy (anonymous):

If you have multiple choice, it's easy enough to just test them all out.

OpenStudy (amistre64):

if you wanna go the long route: y = 2+ax+bx(x+2) y = 2-8x+bx(x+2) y = 2-8x+3x(x+2) and simplify

OpenStudy (anonymous):

Im confused . . .

OpenStudy (anonymous):

Thanks! I was just thinking about that too... testing multiple choice answers is efficient in terms of getting through a hw or test, but if you don't have multiple choice, it's nice to know how to set it up.

OpenStudy (amistre64):

if you are confused, youshould try to lets us know what it is that you are confused about.

OpenStudy (anonymous):

I don't understand how to test them out ?

OpenStudy (anonymous):

The points given are (x,y) coordinates that satisfy the equation. For example, the point (-2,18) means that 18 = a(-2)^2 +b(-2) +c when you put x=-2 and y=18 into the equation, y=ax^2 +bx +c.

OpenStudy (amistre64):

you are given 3 points; which correspond to x and y values: (x,y)

OpenStudy (amistre64):

try equation 1, using (0,2) \[y = –2x^2 – 2x – 3\]replace x with 0 and y with 2 \[2 = –2*0^2 – 2*0 – 3\]and see if the results are true \[2 = – 3~:~false\] therefore it is not equation 1

OpenStudy (anonymous):

If you didn't have multiple choice options, you would have to put all three points into ax^2 + bx +c and solve a system of three equations for a, b, and c.

OpenStudy (amistre64):

... "have to" ? nah. but that would be a suitable method ;)

OpenStudy (anonymous):

so when i use equation 2 it would be like this? y = –3x^2 + 2x – 2 4=-3*42^-3*42-2 ?

OpenStudy (amistre64):

you have the right idea, but it loks like you are using the point (4,42). so youve just replaced the wrong parts

OpenStudy (amistre64):

things are simpler when x=0, so stick with that to narrow the options

OpenStudy (anonymous):

So , I still dont understand how to figure out the answer ?

OpenStudy (amistre64):

y = –3 x^2 + 2 x – 2 ; using (4,42) would be 42=-3*4^2 +2*4 - 2

OpenStudy (amistre64):

but using the (0,2) is way simpler since anyting time 0 is 0 y = –3 x^2 + 2 x – 2 ; using (0,2) would be 2=-3*0^2 +2*0 - 2 2 = -2 since 2=-2 is false, it aint equation 2

OpenStudy (anonymous):

Ohhh , I see . . Hold on .

OpenStudy (anonymous):

y = 3x^2 – 2x + 2 2= 3*0^2-2*0+2 2= 2 Is this the answer , cause I think it is !

OpenStudy (amistre64):

this is a "possible" answer and so is the last equation. so we would need to use one of the other points to verify which of them is correct

OpenStudy (amistre64):

using (4, 42) y = 3 x^2 – 2 x + 2 42 = 3*4^2 – 2*4 + 2 42 = 3*16 – 8 + 2 42 = 48 – 8 + 2 ; i do like this 3rd one, but lets try this point in the last equation to be sure

OpenStudy (amistre64):

using the same point as the last time: (4, 42) y = –2 x^2 + 3 x + 2 42 = –2*4^2 + 3*4 + 2 42 = –2*16 + 12 + 2 42 = –32 + 12 + 2 ; yeah, that point wont work in equation 4; so equation 3 looks great

OpenStudy (anonymous):

Thank you sooo much for helping me , I get it now ! (:

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

Thanks (;

OpenStudy (amistre64):

if that last point DID work in equation 4, then we would have had to use the only point left, and that one would have only fit equation 3

OpenStudy (anonymous):

Yes , I see ! (=

OpenStudy (anonymous):

Enter the points in your stat plot edit screen then push stat, -> calc, quadreg and it gives you the equation.